which fraction is a multiple of 1/2
2/9.....3/2....1/4.... or 1/16

Answer by addingup(3677)

Take 3/2, another way to write it is 1/2+1/2+1/2. Now you should know that 1/2+1/2 = 2/2 = 1, correct? You do know that two halves equals one? So another way to write it is 1 1/2 where 2 halves make the 1 and you have an extra half.